40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 25 10:48:13 2008, Koji, Update, General, Abs. Len. Meas. ~ Realignment / beating / PLL trial 
    Reply  Mon Jul 28 15:54:04 2008, Koji, Update, General, Abs. Len. Meas. ~ More on the beat / the PLL setup PLL_setup.pngbeat_at_8MHz.jpg
       Reply  Thu Jul 31 11:19:30 2008, Koji, Update, General, Abs. Len. Meas. ~ Resonance search trial 
          Reply  Sat Aug 2 12:53:43 2008, Koji, Update, General, Abs. Len. Meas. ~ New PLL at the PSL table yarm_dc.png
             Reply  Mon Aug 4 00:37:58 2008, Koji, Update, General, Abs. Len. Meas. ~ RF PD at the Y end / Manual frequency scan YEND_LAYOUT.png4th_FSR1.png4th_FSR2.png
                Reply  Mon Aug 4 05:23:57 2008, Koji, Update, General, Abs. Len. Meas. ~ Measurement for Y-arm completed 
                Reply  Mon Aug 4 21:48:24 2008, Koji, Update, General, Abs. Len. Meas. ~ Scan for TEM00/01/10 TEM00.pngTEM01.pngknife_edge.png
                   Reply  Mon Aug 4 22:31:10 2008, Koji, Update, General, Abs. Len. Meas. ~ Simple Test for TEM01/10 split TEM01_10.png
                   Reply  Tue Aug 5 00:05:57 2008, Koji, Update, General, Abs. Len. Meas. ~ IFR2023A calibration 
                      Reply  Wed Aug 6 11:10:34 2008, Koji, Update, General, Abs. Len. Meas. ~ analysis of the TEM00 scan TEM00fit.pngTEM00FSRfit.png
                         Reply  Wed Aug 6 11:43:52 2008, Koji, Update, General, Abs. Len. Meas. ~ analysis of the TEM01 scan TEM01fit.png
                            Reply  Thu Aug 14 21:52:51 2008, Koji, Summary, General, Abs. Len. Meas. ~ summary of my Summer mode_spacing_measurement_080816_v2.pdf
Message ID: 793     Entry time: Mon Aug 4 21:48:24 2008     In reply to: 787     Reply to this: 794   795
Author: Koji 
Type: Update 
Category: General 
Subject: Abs. Len. Meas. ~ Scan for TEM00/01/10 
Work log on August 3rd - Part2

o I tried to measure the frequency of the FSRs using TEM00 resonances. Also search of TEM01/TEM10 resonances were tried.

-----------
Measurement for TEM00

o The frequency of the injection beam was scanned from 2MHz to 20MHz using the LabVIEW panel with GPIB. The 1st figure attached below is the result of the scan. Equispaced peaks were found as expected. The interval of the peaks are about 3.89MHz. Each peaks were measured with freq intervals of down to 50Hz. I will analyze the center frequency of the peaks precisely later in order to have a final result.

Measurements for TEM01/TEM10

o The beam injection technique is thought to be useful for measureing the frequency of the higher-order resonances. In order to measure the higher-order resonances the modifications of the experimental setup were applied as below.

1) For TEM10 (the beam like "OO" shape), a razor blade which blocked the horizontal half of the transmitted beam was placed. We needed to disturb half of the beam because the beat between the PSL TEM00 and the injection TEM01 cancels if the PD receives all of the light.

2) The injection beam is slightly misaligned in the horizontal direction in order to enhance the coupling of the injection beam to the cavity TEM01 mode.

3) For TEM01 (the beam like "8" shape), a razor blade cutting the vertical half and the misalignment of the inj beam in the vertical direction are applied.

o The frequency of the injection beam was scaned from 1st FSR of TEM00 in the upward direction. The alignment of the arm cavity was left untouched during the measurement. As shown in the 2nd figure attached below, the resonances were found about 1.19MHz away from the TEM00, but they are separated by about 19kHz(!). This could be split of the degenerated modes which corresponds to the difference of the mirror curvature in two directions! This difference is something like 56 m and 57 m. Can you believe this?

(To be continued to the next entry)
Attachment 1: TEM00.png  78 kB  Uploaded Mon Aug 4 22:50:19 2008  | Hide | Hide all
TEM00.png
Attachment 2: TEM01.png  76 kB  Uploaded Mon Aug 4 23:16:07 2008  | Hide | Hide all
TEM01.png
Attachment 3: knife_edge.png  493 kB  Uploaded Tue Aug 5 03:58:57 2008  | Hide | Hide all
knife_edge.png
ELOG V3.1.3-